Find the equation of the line with a slope of 2.3 and travels through the point (4, 6).
* So first I tried to figure out what is m? what is b? Then try to plug them into the formula: y=mx+b. As a result, I found out that m= 2/3, x=4 and y=-6. The only thing we don't know is what is b.
*-6= 2/3(4)+ b
-6= 2.6666+ b (here is where I got confused because of what I got)
-2.6666=-2.6666 (subtract both sides of the equation by the 2.6666)
b=-8.6666
So then I go to plug in the b to the y=mx+b
-6= 2/3(4)+ (-8.6666)
*It's all supposed to equal -6, but what I get instead is -5.999933333.
Am I doing something wrong?

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Find the equation of the line with a slope of 2.3 and travels through the point (4, 6).
* So first I tried to figure out what is m? what is b? Then try to plug them into the formula: y=mx+b. As a result, I found out that m= 2/3, x=4 and y=-6. The only thing we don't know is what is b.
=========
Assuming you mean m = 2/3 and not 2.3:
------
Use y-y1 = m*(x-x1) where (x1,y1) is the point.
y-6 = (2/3)*(x-4)

Answer by Edwin McCravy(20086)

Find the equation of the line with a slope of 2.3 and travels through the point (4, 6).
Plug in m=2.3 x=4, y=6 in

   y = mx+b
   6 = 2.3(4) + b
   6 = 9.2 + b
-3.2 = b

Substitute m=2.3, b=-3.2, x=x, y=y in

   y = mx+b
   y = 2.3x+(-3.2)
   y = 2.3x-3.2

* So first I tried to figure out what is m?
You are given m, the slope, as 2.3

what is b?
You find that as I did above, as b=-3.2

Then try to plug them into the formula: y=mx+b. As a result, I found out that m= 2/3, x=4 and y=-6.
Oh, I see.  You made a typo and typed 2.3 when you meant to type 2/3.
OK.
Plug in m=2/3 x=4, y=6 in

   y = mx+b
   6 = (2/3)(4) + b
   6 = 8/3 + b
Multiply through by 3 to clear the fraction
  18 = 8 + 3b
  10 = 3b
10/3 = b  

Substitute m=2/3, b=10/3, x=x, y=y in

   y = mx+b
   y = (2/3)x+(10/3)

The only thing we don't know is what is b.
*-6= 2/3(4)+ b
Why did you substitute -6 for y? You should have substituted 6
because you said above that it travels through (4, 6), not 
(4, -6).  Did you make another typo on the point?
OK. I'll assume you meant m=2/3 and travels through (4, -6), not (4,6).

Plug in m=2/3 x=4, y=-6 in

    y = mx+b
   -6 = (2/3)(4) + b
   -6 = 8/3 + b
Multiply through by 3 to clear the fraction
  -18 = 8 + 3b
  -26 = 3b
-26/3 = b  

Substitute m=2/3, b=-26/3, x=x, y=y in

   y = mx+b
   y = (2/3)x+(-26/3)
   y = (2/3)x-26/3

-6= 2.6666+ b (here is where I got confused because of what I got)
-2.6666=-2.6666 (subtract both sides of the equation by the 2.6666)
b=-8.6666
That's approximately correct.  However you should be using
fractions, not rounded off decimals.  -26/3 is only approximately
equal to -8.6666, and BTW if you were going to round off, it would
have been -8.6667, not -8.6666.  But you should have used the 
fraction -26/3.

So then I go to plug in the b to the y=mx+b
-6= 2/3(4)+ (-8.6666)

No. After you have found the slope, you don't substitute the point
again as you did to find b.  The second time you substitute you leave 
x as just the letter x, and y as just the letter y.  You don't 
substitute anything for them to get the final equation.

So the equation should be

y = (2/3)x + (-26/3)

*It's all supposed to equal -6, but what I get instead is -5.999933333.
Am I doing something wrong?
Yes, you are using a rounded off decimal instead of the fraction. On
top of that, you rounded off incorrectly, and made two typos, once
in typing 2.3 for 2/3 and again typing (4,6) for (4,-6).

Edwin